# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ 2.4 Framework for Down-Monotone Problems
  • the total-variation distance $\mathtt{TV}(\mathcal{E},\mathcal{D})<k\sqrt{\frac{\log(2k/\delta)}{2m}}$.
@@
Consider a stochastic problem that is down-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{E}\preceq_{\mathsf{SD}}\mathbf{D}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$, and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
@@
The proof of this lemma is identical to that in §2.2 for up-monotonicity. Using Theorem 2.4 and Lemma 2.5, the proof of Theorem 1.1 (for down-monotone problems) is then identical to that for up-monotonicity (done earlier).
@@ 2.5 Framework for Maximization Problems
Consider a stochastic maximization problem that is down-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\mathtt{TV}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$ and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
@@
Lemma 2.6 then implies Theorem 1.1 for maximization, down-monotone problems: the proof is identical to that for minimization problems (see §2.1). Finally, the same guarantee holds when $\mathcal{P}$ is up-monotone (per the discussion in §2.4).
2.6 Framework for Continuous Distributions
The algorithmic framework described above assumed that the random variables are discrete. Here, we show how to handle stochastic problems with arbitrary (continuous or discrete) r.v.s as long as the policies are based on a finite number of thresholds. We first define a natural discretization of continuous r.v.s.
Definition 2.1 (Thresholded r.v.).
For any $[0,U]$-bounded r.v. $\mathsf{X}$ (possibly continuous) and list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of increasing threshold values, let $\theta(\mathsf{X},\mathbf{{b}})$ be the r.v. such that
$\theta(\mathsf{X},\mathbf{{b}})=i\text{ if }b_{i-1}\leq\mathsf{X}<b_{i},\qquad\forall i\in[k],$
where we use $b_{0}:=0$ and $b_{k}:=U$. This corresponds to thresholding $\mathsf{X}$ according to $\mathbf{{b}}$.
Note that $\theta(\mathsf{X},\mathbf{{b}})$ is a discrete random variable with support size $k$, even if $\mathsf{X}$ is continuous. We will assume that the $\alpha$-approximation algorithm $\mathtt{ALG}$ for the stochastic problem $\mathcal{P}$ always finds a $k$-threshold policy, defined as follows. Such a policy is given by a decision tree $\sigma$, where each node is labeled by an item $j$ to probe and list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of thresholds, such that branch $i\in\{1,\cdots k\}$ is followed when $\mathsf{X}_{j}\in[b_{i-1},b_{i})$. This limits the number of branches out of each node to at most $k$, whereas an arbitrary policy may have an unbounded number of branches (corresponding to each possible outcome). We emphasize that the thresholds at different nodes may be completely different. This restriction allows us to analyze policies involving continuous r.v.s in the same manner as before (for discrete r.v.s). In particular, all other definitions related to the policy (terminal/leaf nodes, policy execution, cost etc) remain the same as for discrete r.v.s. Moreover, for problems like Prophet inequality and Pandora’s Box, there are optimal $1$-threshold policies: so this restriction is without loss of generality for these problems.
Below, we focus on problems involving continuous r.v.s. that are up-monotone and have a minimization objective. We note that this approach also works directly for down-monotone problems or maximization objectives. We need to modify the method to compute stochastically dominating distributions and the stability lemma: these changes are described below.
@@ 2.6 Framework for Continuous Distributions
There is an algorithm that, given $m$ i.i.d. samples from an any bounded distribution $\mathcal{D}$ and parameter $\delta>0$, computes a (discrete) distribution $\mathcal{E}$ that satisfies the following properties with probability at least $1-\delta$:
  • $\mathcal{E}$ is stochastically dominates $\mathcal{D}$.
  • For any list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of $k-1$ thresholds, we have
    $\mathtt{TV}(\theta(\mathsf{X},\mathbf{{b}}),\theta(\mathsf{Y},\mathbf{{b}}))<k\sqrt{\frac{\log(2(k+m)/\delta)}{2m}},\,\,$
    where $\mathsf{X}$ (resp. $\mathsf{Y}$) is distributed according to $\mathcal{D}$ (resp. $\mathcal{E}$).
Proof Outline.
The algorithm is a natural extension of that for the discrete case, and is described in Algorithm 3. The key idea is to put an $\epsilon$ mass on the highest value $U$ and remove an equal mass from the lowest sampled values.
Algorithm 3 StocDomCont
@@
1:Input: samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$, range $[0,U)$, confidence parameter $\delta$
2:let $\epsilon=\min\left\{\sqrt{\frac{\log(2(k+m)/\delta)}{2m}},1\right\}$
3:let $a_{1}\leq\cdots\leq a_{m}$ denote the sampled values, and $a_{m+1}:=U$
4:let $\widehat{p}_{i}=\frac{1}{m}$ for $i\in[m]$ and $\widehat{p}_{m+1}=0$ denote the sample distribution
5:let $y$ be the minimum index such that $\sum_{i=1}^{y}\widehat{p}_{i}\geq\epsilon$, and set
$\bar{p}_{i}=\begin{cases}\widehat{p}_{i}+\epsilon&\text{if }i=m+1\\ \widehat{p}_{i}&\text{if }y<i\leq m\\ \sum_{j=1}^{y}\widehat{p}_{j}-\epsilon&\text{if }i=y\\ 0&\text{if }i<y\end{cases}$ (23)
6:Output: distribution $\mathcal{E}$ with support $\{a_{i}\}_{i=1}^{m+1}$ and probabilities $\{\bar{p}_{i}\}_{i=1}^{m+1}$
It is easy to see that $\mathcal{E}$ is a valid distribution: note that $\bar{p}\geq 0$ and $\sum_{i=1}^{m+1}\bar{p}_{i}=1$. We are now ready to prove the two properties. Consider any list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of thresholds for the second property. Let $c_{1}\leq\cdots\leq c_{r}$ be the values in $\{a_{i}\}_{i=1}^{m+1}\cup\{b_{j}\}_{j=1}^{k-1}$ in increasing order. Note that the $c_{i}$s contain both the sampled values and the thresholds, and $r\leq m+k$. Moreover, the last value $c_{r}=a_{m+1}=U$. We extend the probabilities in $\widehat{p}$ and $\bar{p}$ to all values $c_{1}\leq\cdots\leq c_{r}$ (just set the probability to zero for non-sampled values). To keep notation simple, we will still refer to these “extended” distributions as $\widehat{p}$ and $\bar{p}$.
We again use the empirical distribution as an intermediate quantity: let $\widehat{\mathcal{E}}$ denote the distribution with support $\{c_{i}\}_{i=1}^{r}$ and probabilities $\{\widehat{p}_{i}\}_{i=1}^{r}$; let $\mathsf{Z}$ denote the corresponding r.v. For any $i\in[r]$, let $q_{i}=\mathbf{P}(\mathsf{X}\geq c_{i})$ and $\widehat{q}_{i}=\mathbf{P}(\mathsf{Z}\geq c_{i})$. Using Hoeffding’s bound and union bound (as before), and the value $\epsilon=\sqrt{\frac{\log(2(m+r)/\delta)}{2m}}$,
$\mathbf{P}\left(\exists i\in[r]\,:\,|\widehat{q}_{i}-q_{i}|\geq\epsilon\right)\leq\delta.$

### v2

@@ 2.3 Framework for Down-Monotone Problems
  • the total-variation distance $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E},\mathcal{D})<k\sqrt{\frac{\log(2/\delta)}{2m}}$.
@@
Consider a stochastic problem that is down-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{E}\preceq_{\mathsf{SD}}\mathbf{D}$ and $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$, and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
@@
The proof of this lemma is identical to that in Section 2.1 for up-monotonicity. Using Theorem 2.4 and Lemma 2.5, the proof of Theorem 1.1 (for down-monotone problems) is then identical to that for up-monotonicity (done earlier).
@@ 2.4 Framework for Maximization Problems
Consider a stochastic maximization problem that is down-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions such that $\mathbf{D}\preceq_{\mathsf{SD}}\mathbf{E}$ and $\operatorname{\mathtt{TV}-\mathtt{Dist}}(\mathcal{E}_{i},\mathcal{D}_{i})\leq\epsilon_{i}$ for each $i\in[n]$. If $\sigma$ is the policy returned by $\mathtt{ALG}(\mathbf{E})$ and $\sigma^{*}$ is an optimal policy under $\mathbf{D}$, then:
@@
Lemma 2.6 then implies Theorem 1.1 for maximization, down-monotone problems: the proof is identical to that for minimization problems Finally, the same guarantee holds when $\mathcal{P}$ is up-monotone (per the discussion in §2.3).
3 Handling General Distributions
In this section, we extend results from the discrete setting to continuous (or even mixed) distributions. Here, each random variable $\mathsf{X}_{i}\sim\mathcal{D}_{i}$ is supported over a bounded interval $[a,b]$.
We will assume that there is an $\alpha$-approximation algorithm $\mathtt{ALG}$ for the stochastic problem $\mathcal{P}$ that always returns a $k$-threshold policy, defined as follows. The policy is represented by a decision tree $\sigma$, where each internal node is labeled by an item $i\in[n]$ to probe, along with a list of thresholds $\langle\tau_{1},\dots,\tau_{k-1}\rangle$ that define a partition of $[a,b]$, into intervals $\{J_{1},\dots,J_{k}\}$. Each branch $s\in[k]$ is followed when the observed value $x_{i}\in J_{s}$. This ensures that each node has at most $k$ outgoing branches. We also assume that for any $k$-threshold policy $\sigma$, the cost function $f(\sigma,\mathbf{x})$ can be expressed in the following form:
$f(\sigma,\mathbf{x})=\sum_{\ell:\text{leaf of }\sigma}\left(\prod_{i\in P_{\ell}}\mathbf{I}_{x_{i}\in J_{\ell}^{i}}\right)f_{\sigma,\ell}(\mathbf{x}),\qquad\forall~\mathbf{x}\in[a,b]^{n},$ (23)
where $f_{\sigma,\ell}(\mathbf{x})$ is the cost function associated with the leaf $\ell$, $P_{\ell}$ is the set of items probed along the path to $\ell$ and $J^{i}_{\ell}$ is the interval corresponding to the branch out of item $i$ leading to leaf $\ell$. We note that, given a realization $\mathbf{x}$, only one summand in (23) will be non-zero (the one corresponding to the leaf $\ell$ that $\mathbf{x}$ traces). In this section, we use $f_{\sigma}(\mathbf{x}):[a,b]^{n}\rightarrow\mathbb{R}_{+}$ to denote the function $f(\sigma,\mathbf{x})$ for a fixed policy $\sigma$.
Assumption on the objective function. We need to make some continuity and boundedness assumptions on the leaf functions $f_{\sigma,\ell}$. In order to state them formally, we start with some definitions.
Definition 3.1 (BV function).
The total variation of a 1-dimensional function $g:[a,b]\rightarrow\mathbb{R}$ is:
$V_{a}^{b}(g)=\sup_{\mathcal{P}}\sum_{i=0}^{m}|g(x_{i})-g(x_{i-1})|.$
where the supremum is taken over all finite partitions ${\cal P}=\langle a=x_{0}<x_{1}<\cdots x_{m}=b\rangle$ of the interval. We say that $g$ is a bounded variation (BV) function if and only if $V_{a}^{b}(g)<\infty$.
We list several useful properties of BV functions (Heil, 2019).
  • Any monotone function $g$ over $[a,b]$ is BV and $V_{a}^{b}(g)=|g(b)-g(a)|$.
  • If $g$ is a Lipchitz function over $[a,b]$ then it is BV and $V_{a}^{b}(g)\leq L(b-a)$, where $L$ is the Lipchitz constant.
  • For any BV function $g$ over $[a,b]$, we have $|g(x)|\,\leq\,|g(a)|\,+\,V_{a}^{b}(g)$ for all $x\in[a,b]$.
We now introduce a generalization of BV to high dimensions. Informally, this requires every 1-dimensional “slice” of $f$ to be BV.
Definition 3.2 (Coordinate-wise BV).
A $n$-dimensional function $f(\mathbf{x}):[a,b]^{n}\rightarrow\mathbb{R}$ is said to have coordinate-wise bounded variation if the following conditions hold.
  • For any index $i\in[n]$ and $\mathbf{x}_{-i}=(x_{1},\dots,x_{i-1},x_{i+1},\dots x_{n})\in[a,b]^{n-1}$, the 1-dimensional function $f(y|\mathbf{x}_{-i}):=f(x_{1},\dots,x_{i-1},y,x_{i+1},\dots,x_{n})$ is BV.
  • Moreover, the total variation $(\mathtt{TV})$ of $f(y|\mathbf{x}_{-i})$ is uniformly bounded for all index $i$ and $\mathbf{x}_{-i}$, i.e., there is a finite value $\mathtt{TV}(f)$ such that
    $\sup_{i,\mathbf{x}_{-i}}V_{a}^{b}(f(y|\mathbf{x}_{-i}))\leq\mathtt{TV}(f).$
Similarly, we define coordinate-wise continuous functions as follows.
Definition 3.3 (Coordinate-wise Continuous).
A $n$-dimensional function $f(\mathbf{x}):[a,b]^{n}\rightarrow\mathbb{R}$ is coordinate-wise continuous if the 1-dimensional function $f(y|\mathbf{x}_{-i}):=f(x_{1},\dots,x_{i-1},y,x_{i+1},\dots,x_{n})$ is continuous for all $i\in[n]$ and $\mathbf{x}_{-i}=(x_{1},\dots,x_{i-1},x_{i+1},\dots x_{n})\in[a,b]^{n-1}$.
We assume that the objective function is non-negative and bounded, with $f(\sigma,\mathbf{x})\leq f_{\max}$. Additionally, we make the following (mild) assumptions on the “leaf functions” $f_{\sigma,\ell}$ for any $k$-threshold policy.
Assumption 3.1.
For any $k$-threshold policy $\sigma$ and leaf $\ell$, we assume that $f_{\sigma,\ell}(\mathbf{x})$ is coordinate-wise BV and coordinate-wise continuous. Furthermore, we assume that the total variation of $f_{\sigma,\ell}$ is uniformly bounded over all $\sigma$ and $\ell$, i.e., there is a finite value $\mathtt{TV}_{\max}(f)$, called the total-variation of $f$, such that
$\sup_{\sigma,\ell}\,\mathtt{TV}(f_{\sigma,\ell})\,\leq\,\mathtt{TV}_{\max}(f).$
Preliminaries on probability distribution functions. We will be working with general distributions (discrete or continuous) here. To this end, we now introduce some definitions.
Definition 3.4 (Cumulative Distribution Function).
Given a probability distribution $P$ over $\mathbb{R}^{n}$, we denote by $F_{P}$ the cumulative distribution function (CDF) of $P$, defined as:
$F_{P}(x_{1}\dots x_{n})=\Pr_{(\mathsf{X}_{1}\dots\mathsf{X}_{n})\sim P}[\mathsf{X}_{1}\leq x_{1},\ldots,\mathsf{X}_{n}\leq x_{n}].$
Moreover, if $P=P_{1}\times\dots\times P_{n}$ where $\{P_{i}\}_{i\in[n]}$ is a product distribution, we have:
$F_{P}(x_{1}\dots x_{n})=\prod_{i=1}^{n}F_{P_{i}}(x_{i}).$
Definition 3.5.
The Kolmogorov–Smirnov(KS) distance between two distributions $P$ and $Q$ on $\mathbb{R}$ is defined as :
$d_{\mathrm{KS}}(P,Q):=\sup_{x\in\mathbb{R}}|F_{P}(x)-F_{Q}(x)|.$
In our proof, we require a few standard results from analysis (stated below).
Theorem 3.1 (Combination of Lemma 5.2.14 and Theorem 5.2.15 from Heil (2019)).
If $f:[a,b]\rightarrow\mathbb{R}$, then the following two statements are equivalent.
  • $f\in\mathrm{BV}[a,b]$.
  • There exist monotone non-decreasing functions $f^{+}$ and $f^{-}$ such that $f=f^{+}-f^{-}$. And $V_{a}^{b}(f)=f^{+}(b)+f^{-}(b)-f^{+}(a)-f^{-}(a)$
Theorem 3.2 (Theorem 12.14 from Gordon (1994)).
Let $f$ and $\phi$ be bounded functions defined on $[a,b]$. If $f$ is Riemann-Stieltjes integrable with respect to $\phi$ on $[a,b]$, then $\phi$ is Riemann-Stieltjes integrable with respect to $f$ on $[a,b]$ and
$\int_{a}^{b}\phi df=f(b)\phi(b)-f(a)\phi(a)-\int_{a}^{b}fd\phi$
Theorem 3.3 (Theorem 12.15 from Gordon (1994)).
Let $f$ and $\phi$ be bounded functions defined on $[a,b]$. If $f$ is continuous on $[a,b]$ and $\phi$ is of bounded variation on $[a,b]$, then $f$ is Riemann-Stieltjes integrable with respect to $\phi$ on $[a,b]$.
The following theorem allows us to bound the difference in expectation of a BV function over two different distributions.
@@ 3 Handling General Distributions
Let $g:[a,b]\rightarrow\mathbb{R}$ be a BV function, and $P$ and $Q$ be probability distributions on $\mathbb{R}$. If the Riemann–Stieltjes integrals $\int_{a}^{b}g(x)\,dF_{P}(x)$ and $\int_{a}^{b}g(x)\,dF_{Q}(x)$ exist, then:
$\left|\int_{a}^{b}g\,dF_{P}-\int_{a}^{b}g\,dF_{Q}\right|\leq\left(2g_{\max}+V_{a}^{b}(g)\right)\cdot d_{\mathrm{KS}}(P,Q).$
where $g_{\max}:=\max_{x\in[a,b]}|g(x)|$.
Proof.
By the assumption, we have:
$\left|\int_{a}^{b}g\,dF_{P}-\int_{a}^{b}g\,dF_{Q}\right|\,\,$ $=\,\,\left|g(b)\left(F_{P}(b)-F_{Q}(b)\right)-g(a)\left(F_{P}(a)-F_{Q}(a)\right)-\int_{a}^{b}(F_{P}-F_{Q})\,dg\right|$
$\leq\,\,\left|g(b)\left(F_{P}(b)-F_{Q}(b)\right)\right|+\left|g(a)\left(F_{P}(a)-F_{Q}(a)\right)\right|+\int_{a}^{b}\left|F_{P}-F_{Q}\right|\,dg$
  $\leq\,\,2g_{\max}\cdot d_{\mathrm{KS}}(P,Q)+\int_{a}^{b}\left|F_{P}-F_{Q}\right|\,dg.$
The equality follows from the integration by parts formula (Theorem 3.2).
Using the fact that $g$ is a BV function, we apply the Jordan decomposition (Theorem 3.1) to write $g=g^{+}-g^{-}$, where $g^{+}$ and $g^{-}$ are non-decreasing functions, and $V_{a}^{b}(g)=g^{+}(b)+g^{-}(b)-g^{+}(a)-g^{-}(a)$. We now have
$\int_{a}^{b}\left|F_{P}(x)-F_{Q}(x)\right|\,dg(x)=\int_{a}^{b}\left|F_{P}(x)-F_{Q}(x)\right|\,dg^{+}(x)-\int_{a}^{b}\left|F_{P}(x)-F_{Q}(x)\right|\,dg^{-}(x)$
$\leq\sup_{x\in\mathbb{R}}\left|F_{P}(x)-F_{Q}(x)\right|\cdot(g^{+}(b)-g^{+}(a))+\sup_{x\in\mathbb{R}}\left|F_{P}(x)-F_{Q}(x)\right|\cdot(g^{-}(b)-g^{-}(a))$
$=V_{a}^{b}(g)\cdot d_{\mathrm{KS}}(P,Q),$
This establishes the theorem. ∎
Online to offline framework.
The overall algorithm and analysis is similar to that in the discrete case. Here, we focus on stochastic problems satisfying Assumption 3.1 and up-monotonicity. We also assume that there is an $\alpha$-approximation algorithm via $k$-threshold policies. These results also extend to down-monotone problems in a straightforward way.
At each time step $t=1,\cdots T$, we first construct a (modified) empirical distribution $\mathbf{E}^{t}$ that stochastically dominates the true distribution $\mathbf{D}$ w.h.p. Then, we obtain a $k$-threshold policy $\sigma^{t}$ by running the $\alpha$-approximation algorithm on distribution $\mathbf{E}^{t}$. Finally, we run policy $\sigma^{t}$ on a realization $\mathbf{x}^{t}\sim\mathbf{D}$ and observe semi-bandit feedback.
In the next two subsections, we prove the sampling lemma (that constructs a stochastically dominating distribution) and the stability lemma (that bounds the single-step regret).
3.1 Sampling Algorithm
Suppose the underlying distribution $\mathcal{D}$ has CDF $F(x)$. We will show how to construct an empirical distribution $\mathcal{E}$ that dominates $\mathcal{D}$ where $\mathcal{D}$ and $\mathcal{E}$ are close enough in terms of KS distance.
Theorem 3.5.
@@ 3.1 Sampling Algorithm
There is an efficient algorithm that, given $m$ i.i.d. samples from a distribution $\mathcal{D}$ with bounded support $[a,b]$, computes a distribution $\mathcal{E}$ satisfying the following properties with probability at least $1-\delta$.
  • $\mathcal{E}$ stochastically dominates $\mathcal{D}$.
  • The KS distance $d_{KS}(\mathcal{E},\mathcal{D})<2\sqrt{\frac{\log(2/\delta)}{2m}}$.
Proof.
The main idea is to first construct the empirical distribution, and subsequently move an $\epsilon$ mass from low values to the maximal value $b$. See Algorithm 3 for a formal description of our construction.
Algorithm 3 EmpStocDomCts
@@
1:Input: support $S:=[a,b]$, samples $x_{1},\ldots,x_{m}\sim\mathcal{D}$, confidence parameter $\delta$
2:let $\epsilon=\sqrt{\frac{\log(2/\delta)}{2m}}$
3:let $F_{m}(x)\leftarrow\frac{1}{m}\cdot\sum_{j=1}^{m}\mathbf{I}[x_{j}\leq x]$
4:let $\widehat{F}_{m}(x)=(F_{m}(x)-\epsilon)^{+}$ for all $x<b$, and set $\widehat{F}_{m}(x)=1$ for all $x\geq b$.
5:Output: distribution $\mathcal{E}$ over support $S$ with CDF $\widehat{F}_{m}(x)$
$\mathcal{E}$ is a valid distribution.
By our construction, $\widehat{F}_{m}(x)$ is a right continuous and non-decreasing function with $\widehat{F}_{m}(x)=0$ for all $x<a$ and $\widehat{F}_{m}(x)=1$ for all $x\geq b$. Hence it is a valid CDF.
Stochastic dominance and $d_{KS}$ bound.
Using the choice of $\epsilon$ in the algorithm, by the Dvoretzky-Kiefer-Wolfowitz (DKW) Inequality (Theorem 1.4), we have:
$\Pr\left(\sup_{x\in\mathbb{R}}\,|F_{m}(x)-F(x)|\,\,\geq\,\,\epsilon\right)\,\,\leq\,\,\delta.$